OpenStudy (anonymous):

Suppose a computer virus begins by infecting 8 computers in the first hour after it is released. Each hour after that, each newly infected computer causes 8 more computers to become infected. The function y = 8x models this situation. Make a table with integer values of x from 1 to 4.

OpenStudy (anonymous):

That is incorrect.  Notice that each of the 8 initially infected computer adds 8 newly infected computers. Since each one does that and there are 8, we have \((8+8+8+8)+(8+8+8+8)=8*8\) newly infected computers For the third hour you will have \(8(8) \) or \(64\) each one adding 8 newly infected computers so now we have \(64\) different \(8's\) one for each already infected computer. So \(f(x)=8^x\), you have an exponential function. Just plug in \(x=\{1,2,3,4\}\) to get your table.